1 Let A [1 0 1 1 2 Find A-1.
Answer #1

Soln liven A 1 01 2 det A = (I-2) -I (o) ICo) det A = 3 Mol = -3 To calculate the Inverse matrix find minols and cofactors of- 2.a C32 -1! ;)=2 (1) 5 M32 lo il = 1143 bevejlet (-1)3 +3 M33 matrix of cofactors Erai - ☺ 2 G3 = 3 -2 3 3 # please give a